Let f(x) = 3x - 1, h(x)=x-4. Find (f o f)(-5). (f of)(-5) =
Answer #1

Sol Glven f(x) = 3-1 ha= x - 4 (fof) (-5) = ? Now (Hof)(x) = f(f(n)) $(34-1) 3(3x-1) - 9x-3-1 (fof)(x) = 91-4 =) (404)(-5) =